youdointlikeme:

Compare the two fractions using <, >, or =. 1/2___ 4/10 > < =

jimthompson5910:

Here's a similar example Let's say we want to compare 1/3 and 7/8 First we assume they are equal 1/3 = 7/8 1*8 = 3*7 ... cross multiply 8 = 21 The third equation is false, so the first equation is false We can fix things by replacing each equal sign with a less than sign 1/3 < 7/8 1*8 < 3*7 8 < 21 So we see that 1/3 is smaller than 7/8 You'll follow similar steps in your question. The answer may or may not be a less than sign.

Imagine:

1/2=5/10th's

Imagine:

5/10 _ 4/10

Imagine:

@youdointlikeme

Imagine:

<=Means that 4/10ths are more than 5/10ths

Imagine:

which makes no sense, so > is correct.